OpenStudy (idealist10):

Solve for y: -sin^-1 y=sin^-1 x+C

OpenStudy (idealist10):

@ganeshie8

OpenStudy (idealist10):

It's sine inverse.

ganeshie8 (ganeshie8):

what exactly do u feel is difficult here ?

OpenStudy (idealist10):

The negative sign in sin^-1 y.

ganeshie8 (ganeshie8):

multiply -1 both sides, it goes away eh ?

OpenStudy (idealist10):

So I got sin^-1 y=-sin^-1 x-C

ganeshie8 (ganeshie8):

\[\large -\sin^{-1} y=\sin^{-1} x+C\] \[\large \sin^{-1} y=-\sin^{-1} x-C\]

ganeshie8 (ganeshie8):

yes, take "sin" both sides

ganeshie8 (ganeshie8):

\[\large -\sin^{-1} y=\sin^{-1} x+C\] \[\large \sin^{-1} y=-\sin^{-1} x-C\] \[\large \sin(\sin^{-1} y)=\sin(-\sin^{-1} x-C)\]

OpenStudy (idealist10):

I got y=-sin(sin^-1 x+C)

ganeshie8 (ganeshie8):

left side, \(\sin\) and \(\sin^{-1}\) eat eachother out leaving you : \[\large y =\sin(-\sin^{-1} x-C)\]

ganeshie8 (ganeshie8):

yes, that looks good since "sin" is an odd function, so you can pull out the negative sign : sin(-x) = -sin(x)

OpenStudy (idealist10):

So if I simplify more, I got y=-x-sin*C=-x-c since sine gets absorbed into C, the constant. Am I right?

ganeshie8 (ganeshie8):

NOPE. sin(A+B) \(\large \color{red}{\ne}\) sinA + sinB

ganeshie8 (ganeshie8):

\[\large y =-\sin(\sin^{-1} x+C)\] is the final simplified form, you're done !

OpenStudy (idealist10):

But I have to use the identity sin(A-B)=sinAcosB-cosAsinB.

OpenStudy (idealist10):

Thanks.